# Nonlinear dynamics of confined cell migration—modeling and inference

### v1

@@ II Model
We model a single cell within the phase field framework [19, 20, 21, 22, 9, 23], which describes objects with arbitrary and continuously deforming shapes. Because the experiments of [14] confine the cell on two-dimensional adhesive micropatterns, we simplify our model to two dimensions and adopt a “top view” of the cell (see Fig. 3). The cell shape is set by a field $\phi({\bm{r}},t)$, which is zero outside the cell and one inside it, implicitly defining the boundary at $\phi=1/2$. The evolution of the cell boundary is governed by energy minimization and advection of the interface [20, 24, 25],
@@
where $\mathcal{M}$ is the transport coefficient, and ${\bm{v}}({\bm{r}},t)$ is the velocity field of the cell driving its interface. The free energy is $\mathcal{F}=\mathcal{F}_{\mathrm{CH}}+\mathcal{F}_{\mathrm{area}}$. The Cahn-Hilliard energy is [20]
@@
The left hand side represents the friction force per unit area, while the first term on the right denotes the velocity of the cell boundary as governed by energy minimization. The last term is the active motility force per unit area, which depends on cell polarity and drives the cell front.
@@ II.1 Micropattern geometries
The experiments of [14] use an adhesive fibronectin micropattern to effectively confine the cells—cells can only adhere to the substrate on the micropattern. We incorporate these micropatterns in our phase field framework using a static field $\chi(\bm{r})$. We set $\chi$ to be zero in the interior and one in the exterior of the micropattern, transitioning smoothly at the boundary (Fig. 1). We assume that the cell interacts with the substrate through its polarity, which controls cell motility. This neglects any adhesive energy to the micropattern [26, 21].
@@ II.2 Defining cell polarity and motility
Here, $\alpha$ is a constant, $\bm{\hat{n}}\equiv-\nabla\phi/|\phi|$ is the unit vector normal to the cell interface, and the phase field terms ensure that motility forces exist only where the cell interface (i.e. $\phi^{2}(1-\phi)^{2}>0$) is in contact with the micropattern (i.e. $01-\chi>$ 0)[28, 21, 33].
@@
The second term in Eq. (6) says that cells decrease their polarity in regions where they lose contact with the micropattern (places where $\chi\phi\neq 0$). This term, along with the filopodial probability in Eq. 7, are the only two mechanisms by which the cell interacts with the micropattern.
The third term in Eq. (6) uniformly inhibits polarity when the cell perimeter $s$ exceeds a threshold $s_{0}$ (indicated by the Heaviside function $H(s-s_{0})$). This assumption is motivated by experiments where membrane tension was observed acting as a global, long-range inhibitor of actin polymerization in neutrophils [37] (also see [38, 39]).
@@ II.3 Parameter setting
Parameter
Value
Units
$\gamma$
0.8
$\gamma_{0}$
$\lambda$
4.8
${\mu\mathrm{m}}$
$\kappa$
6000
$\gamma_{0}\ {\mu\mathrm{m}}$
$R_{0}$
18
${\mu\mathrm{m}}$
$\eta$
0.67
$\gamma_{0}\ \rm min\ {\mu\mathrm{m}}^{-3}$
$\mathcal{M}$
0.75
${\mu\mathrm{m}}\ (\gamma_{0}\ \rm min)^{-1}$
$\mu_{\beta}$
0.94
$\rm min^{-1}$
$\sigma_{\beta}$
4.7
$\rm min^{-1}$
$\sigma$
6.7
${\mu\mathrm{m}}$
$s_{0}$
$2\pi(1.5R_{0})$
${\mu\mathrm{m}}$
$\tau_{f}$
2.2
$\rm s$
$\tau_{\chi}$
9.6
$\rm s$
$\tau_{s}$
8
$\rm min$
$\tau$
4
$\rm min$
$\ell$
6
${\mu\mathrm{m}}$
$\alpha$
1.4
$\gamma_{0}\ {\mu\mathrm{m}}^{-2}$
@@
The constraint penalizing deviations from the preferred area, $\kappa$, is set relatively strongly to ensure the area of the default cell varies less than $0.5\%$ from its target. The timescale over which cell polarity decays to zero in regions where the cell is near substrate edges, $\tau_{\chi}$, is set so that when the cell boundary leaves the micropattern, the polarity rapidly becomes negative, leading to the membrane contracting away from the micropattern edge, preventing the cell from leaving the pattern entirely. The perimeter-induced degradation in polarity prevents the cell from becoming too elongated. Since we do observe that cells in [14] can be quite elongated, we cannot set the timescale of this process $\tau_{s}$ too short, or it will prevent cell stretching. We choose $\tau_{s}$ to be on the order of the global decay $\tau$, and note that the decay rate will be faster for larger values of cell perimeter. We set the magnitude of the motility force $\alpha$ so that our simulated cells achieve speeds and accelerations that roughly match the experiments. The rest of the parameters in Table 1 will be discussed in §III as we consider their variations and effect on cell behavior.
@@ II.4 Inferring motility patterns from simulations
We begin each simulation by either positioning the cell randomly in one of the two basins of the two-state substrate or uniformly distributing it across the rectangular micropattern. Upon initialization, we numerically evolve the phase field equations of motion for 48 hours (see Appendix A). Unlike the experiments, which recorded the position of cell nucleus, we track the $x$ position of cell center of mass. While nuclear dynamics could be modeled with another phase field [9], we neglect it here for simplicity and claim that cell center of mass is a fair approximation of cell nucleus for most of the morphologies we observe.
@@
where $\eta(t)$ is a Gaussian white noise with $\langle\eta(t)\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, we focus only on $F(x,v)$, which denotes the average center-of-mass acceleration of the cell when it has position $x$ and velocity $v$. Since cell motility is stochastic, we perform $N=4800$ simulations for every cell-micropattern pair. Mirroring the approach of [14], we compute velocities and accelerations as numerical derivatives, construct the stochastic acceleration field $\dot{v}(x,v)$ from each simulation, and define the deterministic footprint by computing the average acceleration $\dot{v}$ conditional on the value of $x$ and $v$, i.e. $F(x,v)\equiv\langle\dot{v}|x,v\rangle$ (refer to Appendix B for details on this inference scheme).
@@
The blueprints $F(x,v)$ are inferred from a finite number of observations, leading to some uncertainty in the values of $F(x,v)$. This can lead to notable uncertainty in some terminal points and separatrices. We use a bootstrapping technique to show the scale of expected variations in our streamline plots, terminal points, and separatrices in Appendix E.
@@ III.1 Model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns
Figure 3: Overview of limit-cycle (3-3), stationary (3-3), and bistable (3-3) behavior. Each case shows sample snapshots of typical morphologies and polarity (red: high, blue: low) (3, 3, 3), outlines center-of-mass trajectories (3, 3, 3), plots $F(x,v)$ inferred from simulations (3, 3, 3), and reprints the experimental footprints of [14] with permission for comparison (3, 3, 3). Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.
@@ III.1.1 Geometry sensing and the origin of limit-cycles
Let’s consider migration in the two-state geometry. When the cell is inside a basin and large enough to fill it up, most of its interface lies near the edges of the confinement except the part facing the bridge. Since we degrade polarity in regions where the cell overlaps with the micropattern (Eq. (6)), and suppress the generation of activity in regions that are near to the end of the micropattern (filopodial probability of Eq. (7)), sustained polarity can only develop in regions facing the bridge. This region further experiences a targeted increase in average cell polarity due to the positive feedback assumption of Eq. (7). The tendency to polarize toward free micropattern space and positive feedback work cooperatively to yield a persistent, singly-polarized cell that migrates through the bridge and into the other basin. Once on the other side, the cell reverses in polarity as the bridge now faces the opposite direction and repeats the cycle above. The result is a persistently polarized cell that limit-cycles on the two-state geometry. Interestingly, we find that the filopodial probability of Eq. (7) is dispensable—the assumption that leaving the micropattern decreases cell polarity is sufficient for reproducing limit-cycles in the two-state micropattern (see Section III.6 for details).
@@ III.1.2 Origin of stationary behavior on rectangular substrates
Let’s now consider the evolution of cell polarity on the rectangular substrate. Suppose polarity is initially constant in space. Since the cell is confined at its top and bottom, sustained polarity can only develop at the sides, yielding elongation in ±$\pm x$. Given the stochastic evolution of cell polarity and our assumption of positive feedback, which favors a one-side-wins-all configuration, we would initially expect the cell to be singly-polarized and have one clear front, just like it is on the two-state substrate. Instead, we observe the average cell shape is bipolar (see Fig. 3 and Movie M2). This bipolarity arises because of a coupling between the cell polarity and the evolution of cell shape. When we numerically solve the polarity model of Eqs. (6) and (7) on a static phase field (see Appendix C), we consistently arrive at singly-polarized cells. What is different about a cell whose boundary dynamically moves? The polarity field $\mathbb{P}(\mathbf{r},t)$ is confined within the body of the cell, and while it responds to motion of the cell boundary, it does not do so instantaneously. When a high-amplitude patch of activity hits the contour ${\bm{\theta}}_{k}$, it increases the polarity in that region and leads to a protrusion of the cell membrane. This protrusion moves ${\bm{\theta}}_{k}$ normally outward and away from its previous location. Because $\mathbb{P}$ does not instantly migrate with the cell boundary, the polarity measured at the same contour point shortly after protrusion is not nearly as high as before—i.e. polarity can be implicitly suppressed by protrusion. We hypothesize that this negative feedback can prevent our positive feedback terms in Eq. (7) from leading to a singly-polarized cell. Our discovery that cell shape dynamics are necessary here is part of a broader realization that complex behaviors can emerge from the coupling of cell polarization to cell shape [41, 24, 42].
@@ III.2 Cell size modulates bistability on two-state substrates
Although the experiments of [14] do not monitor cell size quantitatively, their example movies show that MDA-MB-231 cells are typically spread over a larger area compared to MCF10A cells. There is also evidence that cancer cells have increased lamellipodial activity compared to their healthy counterparts [43]. Additionally, the absence of the tumor-suppressing micro RNA miR-200c has been shown to result in increased sizes of gastric cancer cells [44]. Collectively, these findings support the notion that cancer cells tend to have larger spread areas, rendering our model’s nomination of cell size as a predictor of motility behavior in two-state confinement plausible.
@@ III.4 Model recapitulates “negative friction” on two-state and normal friction on rectangular geometries
This negative friction is entirely due to the narrow constriction in the two-state micropattern. For cells migrating on rectangular micropatterns or fully constriction-free substrates, we observe a nearly linear acceleration-velocity relationship, which is also consistent with the experiments of [14] (Fig. 5).
@@ III.5 Perturbing the polarization patch model
Figure 6: An overview of how various perturbations in the spatially-correlated activity patch we add to cell polarity affect the deterministic migration of the default cell on the two-state (left column) and rectangular (right column) micropatterns. In each column, the middle figure illustrates the behavior of the default cell, except in the last row, where the left figure represents this. The perturbations considered are in 6 the mean and 6 standard deviation of the magnitude of the noisy activity patch, 6 the spatial extent of the correlated activity patch, and 6 the frequency of adding noisy activity to cell polarity. Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.

### v2

# Inferring nonlinear dynamics of cell migration

@@ II Model
We model a single cell within the phase field framework [19, 20, 21, 22, 9, 23], which describes objects with arbitrary and continuously deforming shapes. Because the experiments of [14] confine the cell on two-dimensional adhesive micropatterns, we simplify our model to two dimensions and adopt a “top view” of the cell (see Fig. 3). The cell shape is set by a field $\phi({\bm{r}},t)$, which smoothly varies from zero outside the cell and one inside it, implicitly defining the boundary at $\phi=1/2$. The evolution of the cell boundary is governed by energy minimization and advection of the interface [20, 24, 25],
@@
where $\mathcal{M}$ is a transport coefficient, and ${\bm{v}}({\bm{r}},t)$ is the velocity field of the cell driving its interface. The free energy is $\mathcal{F}=\mathcal{F}_{\mathrm{CH}}+\mathcal{F}_{\mathrm{area}}$. The Cahn-Hilliard energy is [20]
@@
The left hand side represents the friction force per unit area, while the first term on the right denotes the force density arising from energy minimization. The last term is the active motility force per unit area, which depends on cell polarity and drives the cell front.
@@ II.1 Micropattern geometries
The experiments of [14] use an adhesive fibronectin micropattern to effectively confine the cells—cells can only adhere to the substrate on the micropattern. We incorporate these micropatterns in our phase field framework using a static field $\chi(\bm{r})$. We set $\chi$ to be zero in the interior and one in the exterior of the micropattern, transitioning smoothly at the boundary (Fig. 1). We assume that the cell interacts with the substrate through its polarity, which controls cell motility. This neglects any adhesive energy to the micropattern [26, 21, 27].
@@ II.2 Defining cell polarity and motility
Here, $\alpha$ is a constant, $\bm{\hat{n}}\equiv-\nabla\phi/|\nabla\phi|$ is the unit vector normal to the cell interface, and the phase field terms ensure that motility forces exist only where the cell interface (i.e. $\phi^{2}(1-\phi)^{2}>0$) is in contact with the micropattern (i.e. $01-\chi>$ 0)[29, 21, 34].
@@
The second term in Eq. (6) says that cells decrease their polarity in regions where they lose contact with the micropattern (places where $\chi\phi\neq 0$). This term, along with the filopodial probability in Eq. 7, are the only two mechanisms by which the cell interacts with the micropattern. This inhibition term reflects a signaling interaction between the cell and the adhesive substrate. Contact between the cell and an adhesive substrate can up-regulate and localize Rac activity via focal adhesions and FAK [37, 38]. We treat the absence of this activation as an effective inhibition – so we would expect the strength of the $\chi\phi$ term to decrease if the adhesivity of the substrate decreases or FAK is inhibited.
The third term in Eq. (6) uniformly inhibits polarity when the cell perimeter $s$ exceeds a threshold $s_{0}$, using the Heaviside function $H(s-s_{0})$. This assumption is motivated by experiments where membrane tension was observed acting as a global, long-range inhibitor of actin polymerization in neutrophils [39] (though also see [40, 41]).
@@
We note that Eq. (6) does not have a term advecting the polarity to match the cell boundary motion. This equation thus assumes that the cell’s polarity does not move significantly with respect to the substrate. We often implicitly think of polarity as localized to the basal membrane, making this an assumption that the membrane is relatively strongly linked to the substrate. This is a useful simplifying assumption, often used in models coupling cell shape and polarity [42, 43, 44]. Generalizations are possible, but require mechanical modeling of the cell interior [10, 45] or assumptions relating interior flow to boundary flow [46].
@@ II.3 Parameter setting
Parameter Value Units
$\gamma$ 0.8 $\gamma_{0}$
$\lambda$ 4.8 ${\mu\mathrm{m}}$
$\kappa$ 6000 $\gamma_{0}\ {\mu\mathrm{m}}$
$R_{0}$ 18 ${\mu\mathrm{m}}$
$\eta$ 0.67 $\gamma_{0}\ \rm min\ {\mu\mathrm{m}}^{-3}$
$\mathcal{M}$ 0.75 ${\mu\mathrm{m}}\ (\gamma_{0}\ \rm min)^{-1}$
$\mu_{\beta}$ 7.5 unitless
$\sigma_{\beta}$ 37.5 unitless
$\sigma$ 6.7 ${\mu\mathrm{m}}$
$s_{0}$ $2\pi(1.5R_{0})$ ${\mu\mathrm{m}}$
$\tau_{f}$ 2.2 $\rm s$
$\tau_{\chi}$ 9.6 $\rm s$
$\tau_{s}$ 8 $\rm min$
$\tau$ 4 $\rm min$
$\ell$ 6 ${\mu\mathrm{m}}$
$\alpha$ 1.4 $\gamma_{0}\ {\mu\mathrm{m}}^{-2}$
@@
Many parameters in the model are set fairly roughly to represent large penalty terms. For instance, the constraint penalizing deviations from the preferred area, $\kappa$, is a large value, which ensures the area of the default cell varies less than $0.5\%$ from its target – we do not think of this as a relevant biological parameter but primarily a simple way to set a constraint. Similarly, the timescale over which cell polarity decays to zero in regions where the cell is near substrate edges, $\tau_{\chi}$, is set so that when the cell boundary leaves the micropattern, the polarity rapidly becomes negative, leading to the membrane contracting away from the micropattern edge, preventing the cell from leaving the pattern entirely. The perimeter-induced degradation in polarity prevents the cell from becoming too elongated. Since we do observe that cells in [14] can be quite elongated, we cannot set the timescale of this process $\tau_{s}$ too short, or it will prevent cell stretching. We choose $\tau_{s}$ to be on the order of the global decay $\tau$. We set the magnitude of the motility force $\alpha$ so that our simulated cells achieve speeds and accelerations that roughly match the experiments. The rest of the parameters in Table 1 will be discussed in §III as we consider their variations and effect on cell behavior.
The phenomenological parameters describing cell polarity dynamics like $\tau_{f}$, the frequency of polarity patches, and the magnitude and variability $\mu_{\beta}$ and $\sigma_{\beta}$ of patch polarity, are largely set to recapitulate the core limit cycle behavior for the default cell. Matching these values to other experiments is not straightforward, but the value of $\tau_{f}=2.2$ s is at least plausible – the Rho GTPase cdc42, a key element of mammalian cell polarity, is pulsatile with a timescale of $\sim$ a few seconds [48]. However, the lifetime of protrusions may be quite different than this value ($\sim$ minutes, on our rough scale of $\tau$), both in experiment (see e.g. [49]) and in our simulation, given the feedback mechanisms we include. Future work to more precisely calibrate the model would require measurements of protrusion statistics on MDA-MB-231 in the absence of a geometric signal.
We also note that, though some of the values shown in Table 1 are not round numbers, the model is not too carefully tuned – these numbers are essentially historical artifacts of the original fitting process, and were set as some percentage of an original value. Other awkward values arise because of scaling between simulation parameters and physical units. We address the sensitivity of the model in Section III.5 below.
@@ II.4 Inferring motility patterns from simulations
We begin each simulation by either positioning the cell randomly in one of the two basins of the two-state substrate or uniformly distributing it across the rectangular micropattern. Upon initialization, we numerically evolve the phase field equations of motion for 48 hours (see Appendix A). Unlike the experiments, which recorded the position of the cell nucleus, we track the $x$ position of the cell center of mass. While nuclear dynamics could be modeled with another phase field [9], we neglect it here for simplicity and claim that cell center of mass is a fair approximation of cell nucleus for most of the morphologies we observe.
@@
where $\eta(t)$ is a Gaussian white noise with $\langle\eta(t)\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. Here, we focus only on $F(x,v)$, which denotes the average center-of-mass acceleration of the cell when it has position $x$ and velocity $v$. Since cell motility is stochastic, we perform $N=4800$ simulations of 48 hours for every cell-micropattern pair. Mirroring the approach of [14], we compute velocities and accelerations as numerical derivatives, construct the stochastic acceleration field $\dot{v}(x,v)$ from each simulation, and define the deterministic footprint by computing the average acceleration $\dot{v}$ conditional on the value of $x$ and $v$, i.e. $F(x,v)\equiv\langle\dot{v}|x,v\rangle$ (refer to Appendix B for details on this inference scheme).
@@
The blueprints $F(x,v)$ are inferred from a finite number of observations, leading to some uncertainty in the values of $F(x,v)$. This can lead to notable uncertainty in some terminal points and separatrices. We use a bootstrapping technique to show the scale of expected variations in our streamline plots, terminal points, and separatrices in Appendix G.
@@ III.1 Model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns
Figure 3: Overview of limit-cycle (3-3), stationary (3-3), and bistable (3-3) behavior. Each case shows sample snapshots separated by time $\Delta_{t}=10\textrm{min}$ of typical morphologies and polarity (red: high, blue: low) (3, 3, 3), outlines center-of-mass trajectories (3, 3, 3), plots $F(x,v)$ inferred from simulations (3, 3, 3), and reprints the experimental footprints of [14] with permission for comparison (3, 3, 3). Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details.
@@ III.1.1 Geometry sensing and the origin of limit-cycles
Let’s consider migration in the two-state geometry. When the cell is inside a basin and large enough to fill it up, most of its interface lies near the edges of the confinement except the part facing the bridge. Since we degrade polarity in regions where the cell escapes the micropattern (Eq. (6)), and suppress the generation of activity in regions that are near to the end of the micropattern (filopodial probability of Eq. (7)), sustained polarity can only develop in regions facing the bridge. This region further experiences a targeted increase in average cell polarity due to the positive feedback assumption of Eq. (7). The tendency to polarize toward free micropattern space and positive feedback work cooperatively to yield a persistent, singly-polarized cell that migrates through the bridge and into the other basin. Once on the other side, the cell reverses in polarity as the bridge now faces the opposite direction and repeats the cycle above. The result is a persistently polarized cell that limit-cycles on the two-state geometry. Interestingly, we find that the filopodial probability of Eq. (7) is dispensable—the assumption that leaving the micropattern decreases cell polarity is sufficient for reproducing limit-cycles in the two-state micropattern (see Section III.6 for details).
@@ III.1.2 Origin of stationary behavior on rectangular substrates
Let’s now consider the evolution of cell polarity on the rectangular substrate. Suppose polarity is initially constant in space. Since the cell is confined at its top and bottom, sustained polarity can only develop at the sides, yielding elongation in ±$\pm x$. Given the stochastic evolution of cell polarity and our assumption of positive feedback, which favors a one-side-wins-all configuration, we would initially expect the cell to be singly-polarized and have one clear front, just like it is on the two-state substrate. Instead, we observe the average cell shape is bipolar (see Fig. 3 and Movie M2). This bipolarity arises because of a coupling between the cell polarity and the evolution of cell shape. When we numerically solve the polarity model of Eqs. (6) and (7) on a static phase field (see Appendix D), we consistently arrive at singly-polarized cells. What is different about a cell whose boundary dynamically moves? The polarity field $\mathbb{P}(\mathbf{r},t)$ is confined within the body of the cell, and while it responds to motion of the cell boundary, it does not do so instantaneously. When a high-amplitude patch of activity hits the contour ${\bm{\theta}}_{k}$, it increases the polarity in that region and leads to a protrusion of the cell membrane. This protrusion moves ${\bm{\theta}}_{k}$ normally outward and away from its previous location. Because $\mathbb{P}$ does not instantly migrate with the cell boundary, the polarity measured at the same contour point shortly after protrusion is not nearly as high as before—i.e. polarity can be implicitly suppressed by protrusion. We hypothesize that this negative feedback, combined with the stochasticity of the protrusion dynamics, can prevent our positive feedback terms in Eq. (7) from leading to a singly-polarized cell. Earlier work has also found that simple Rho GTPase polarity models can have coexisting peaks over relevant timescales, though there is a tendency of peak merger, but the timescale of coexistence is strongly parameter-dependent [50]. It may thus be possible to reproduce this coexistence by tuning parameters even in a stationary cell. Nevertheless, our discovery that cell shape dynamics can qualitatively change polarization dynamics is part of a broader realization that complex behaviors can emerge from the coupling of cell polarization to cell shape [51, 24, 52].
@@ III.2 Cell size modulates bistability on two-state substrates
Although the experiments of [14] do not monitor cell size quantitatively, their example movies show that MDA-MB-231 cells are typically spread over a larger area compared to MCF10A cells. There is also evidence that cancer cells have increased lamellipodial activity compared to their healthy counterparts [53]. Additionally, the absence of the tumor-suppressing micro RNA family miR-200 has been shown to result in increased sizes of gastric cancer cells [54]. Collectively, these findings support the notion that cancer cells tend to have larger spread areas, rendering our model’s nomination of cell size as a predictor of motility behavior in two-state confinement plausible.
@@ III.4 Model recapitulates “negative friction” on two-state and normal friction on rectangular geometries
This negative friction is entirely due to the narrow constriction in the two-state micropattern. For cells migrating on rectangular micropatterns, we observe a nearly linear acceleration-velocity relationship, which is also consistent with the experiments of [14] (Fig. 5). We also see a similar near-linear behavior in $F(x\to 0,v)$ in unconstrained cells on two-dimensional substrates fully coated with fibronectin (“free 2D” curve of Fig. 5). Our observation that the rectangular and two-dimensional geometries have nearly-identical, nearly-linear curves is also consistent with the experimental results of [14]. A movie of a cell in an unconfined geometry is shown in Movie 5.
@@ III.5 Perturbing the polarization patch model
Figure 6: An overview of how various perturbations in the spatially-correlated activity patch we add to cell polarity affect the deterministic migration of the default cell on the two-state (left column) and rectangular (right column) micropatterns. In each column, the middle figure illustrates the behavior of the default cell, except in the last row, where the left figure ($\tau_{f}=2.2$ s) is the default cell. The perturbations considered are in 6 the mean and 6 standard deviation of the magnitude of the noisy activity patch, 6 the spatial extent of the correlated activity patch, and 6 the frequency of adding noisy activity to cell polarity. Streamlines are colored by equilibrium point; $\times$ indicates equilibrium points where streamlines terminate. Solid black lines are separatrices—crossing this line changes the long-run behavior of the cell. See Section II.4 for details. Results are obtained from 2400 simulations.